57. Find the total derivative dz/dt, given

(a) z = x^2− 8xy − y^3 , where x = 3t and y = 1 − t.

(b) z = f(x, y, t), where x = a + bt, and y = c + k
Answer #1

a) z = x^2− 8xy − y^3 , where x = 3t and y = 1 − t

The formula for total derivative is given by,

dz / dt = zx*dx/dt + zy*dy/dt, where zx = dz / dx and zy = dz / dy.

So,

zx = 2x - 8y and zy = - 8x - 3y2

dz/dt = (2x - 8y)*(3) + ( - 8x - 3y2)*(-1) = (2x - 8y)*(3) + (8x + 3y2)*(1)

dz / dt = (2*3t - 8(1 - t))*(3) + (8*3 - 3(1 - t)2)

dz/dt = (6t - 8 + 8t)*3 + 24 - 3(1 - 2t + t2)

dz/dt = 42t - 24 + 24 - 3 + 6t - 3t2

dz/dt = 48t - 3t2 - 3

---------------------------------------------------------------------------------------------------------------------------------------------------------

b) z = f(x, y, t), where x = a + bt, and y = c + k

dz / dt = (dz / dx)(dx/dt) + (dz/dy)(dy/dt) + (dz/dt)(dt/dt)

dz / dt = zx(b) + zy(0) + zt(1)

dz/dt = b*zx + zt